In a class,the ratio of boys to the girls is 7to8 Which part of the whole class is girls?
Answer by stanbon(75887)

In a class,the ratio of boys to the girls is 7to8 Which part of the whole class is girls?
# of boys = 7x
# of girls = 8x
----
Percent of girls:: 8x/(15x) = 8/15 = 53 1/3%
------------
Cheers,
Stan H.